OpenStudy (anonymous):

How do I find the solution set of |2x – 3| – 10 = –1?

OpenStudy (anonymous):

First isolate your absolute value part.

OpenStudy (anonymous):

Like so: |2x - 3| - 10 = -1 |2x - 3| = 9 From here you do both signs for the right side of the equation. Positive and negative. 2x - 3 = 9 2x = 12 x = 6 2x - 3 = -9 2x = -6 x = -3 So the values that are valid for x are: -3 and 6

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (anonymous):

Np